An 8 percent per year coupon bond with semiannual coupons and $10,000 face value promises to...

  1. An 8 percent per year coupon bond with semiannual coupons and $10,000 face value promises to pay:

A. $800 per year and $10,000 on maturity

B. $80 per year and $1,000 on maturity

C.  $40 every six months and $1,000 on maturity

D. $400 every six months and $10,000 on maturity

  1. 2. The Wall Street Journal shows the bid and ask quotes for a T-bill maturing in 60 days as 0.098 and 0.088, respectively. If you want to sell a 60-day T-bill with face value of $10,000, you will receive (ignore brokerage commissions)

A. $9,836.67

B. $9,998.53

C.  $9,998.37

D. $9,998.39
Answer #1

As per HomeworkLib guidelines when there are more than one question then we have to answer first question.

coupon rate= 0.08/2=0.04

Coupon payment= 10000*0.04= $400

Since it is semiannual so coupon payment will be made every six months.

So correct answer is D) $400 every six months and $10000 on maturity